OpenStudy (anonymous):

A full tank of water springs a leak and empties in three hours. Water escapes the tank at a rate of g(t) gallons per hour, where t, 0

OpenStudy (dumbcow):

use slope definition here \[g'(1.5) = \frac{g(t+h) -g(t)}{h}\] based on available data points use h=.5 \[g'(1.5) = \frac{g(2) -g(1.5)}{.5} = -4\]

OpenStudy (anonymous):

Part (b) is an application of the intermediate value theorem, I believe.